OpenStudy (anonymous):

Divide z^2-9z+14 by z-7 plz show me how to do this

OpenStudy (anonymous):

try starting by factoring the top: \[\frac{ z^2-9z+14 }{ z-7 } = \frac{ (z-7)(z-2) }{ (z-7) }\]

OpenStudy (anonymous):

does that help get you started?

OpenStudy (anonymous):

ok, i'll try

OpenStudy (anonymous):



OpenStudy (anonymous):

??

OpenStudy (anonymous):

use long division. the answer is z-2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

z-2 is correct. I am not sure about the long division technique, but in this example, since you could factor and cancel the (z-7) terms, it allowed you to get z-2... that is correct here. @arisauotani might be able to elaborate on long division technique...

OpenStudy (anonymous):

what about the 9z+14 ?

OpenStudy (anonymous):

my webassign shows z-2 to be incorrect ??

OpenStudy (anonymous):

z^2 -9z + 14 is equal to (z-2)(z-7)... if you multiply out those parenthesis terms, you will get z^2 - 2z - 7z + 14, and the two middle terms simplify to - 9z So on the top of the fraction you have (z-2)(z-7) and on the bottom, you have (z-7) so when you cancel, the (z-7) terms go away and leave (z-2) alone on top. I can't imagine why your website shows z-2 to be incorrect. Is there any more information given in the problem or the directions.

OpenStudy (anonymous):

Divide z2 − 9z + 14by z − 7.

OpenStudy (anonymous):

z^2

OpenStudy (anonymous):

thats the question

OpenStudy (anonymous):

OMG....Im soo SORRY i typed in a capital z, thats why it didnt take it. it took it now. Thank u soo much

OpenStudy (anonymous):

that's so so so annoying, isn't it!!

OpenStudy (anonymous):

glad you figured it out though :) Good luck!

OpenStudy (anonymous):

thank u